Q 5-8 Your company produces electronic connectors, which consist of a metal strip with multiple prongs assembled into a plastic housing. The manufacturing process is divided into four separate departments: 1) Injection molding, where the small plastic housings are produced; 2) Stamping, where the steel strips are punched out of thin coils of steel; 3) Plating, where the steel strips, after being stamped, are plated with tin or gold; and 4) Assembly, where the plated strips are inserted into the plastic housing.

The injection molding machines together have the capacity to produce 4000 plastic housing per hour. The firm's stamping presses can punch out 3000 steel strips per hour. The plating department has two dedicated lines—one for tin-plating and the other for the slower gold-plating process. The tin-plating line has a capacity of 2000 units per hour; the gold-plating line has a capacity of 500 units per hour. In the assembly department, there are six automated assembly machines, each with the capacity to assemble 400 connectors per hour.

Assume that machines never break down, and assume that there is sufficient plastic and steel raw material so that raw material availability never constrains the process. Assume that the plant is only running the tin-plated product—the gold-plating line is completely idle.

Q1: Assume that hourly demand during work hours for tin-plated connectors is 2200 units/hr and demand for gold-plated connectors is 600 units/hr, and that both processes are active. The plant runs for 8 hours/day, 5 days/week, 50 weeks/yr. Both connector types earn a margin of $0.10 apiece. The plant is considering investing in an additional automated assembly machine, which costs $400,000. What is the payback period for this investment (in years)?
Answer #1

Answer: Payback period = 5 years.

5 years will be the pay back period for investing $400000/- in getting a new automated assembly machine which will increase the connector capacity by 400 units/hour.

The current capacity is = capacity of unit machine * total machines = 400*6 = 2400 nos.

The additional machine will increase the capacity by 400 units. So total capacity will be = 2400+400=2800 nos.

The payback period is calculated as below

New Machine Investment Increase Units /hr Daily Capacity Increase @ 8 hours/day Weekly Capacity Increase @ 5 days/week Yearly Capacity Increase @ 50 weeks/Year Margin per piece Total Yearly Margin Increase Investment Payback Period = Investment / yearly increased margin
Capacity increase by additional automated assembly machine 400 3200 16000 800000 $ 0.1 $ 80,000 $400,000 $ 400000 / $80000 = 5 Years